Hello, and good part of day! Please help me!
A pollution control device on a car reduces the amount of an air pollutant by 55 ppm. When the drived is installed, it will take a car 7 hours to emit the same amount of the pollutant that it formerly did in 2h. What was the original concentration of the pollutant in the car's exhaust?
Thank you!

Answer by mananth(16949)

let the original pollutant be x ppm / hour
..
the device reduces by 55 ppm
..
so the pollutant released will be x-55./ hour
..
2x= 7(x-55)
2x=7x-385
add -7x to both sides
2x-7x=7x-7x-385
-5x=-385
divide by -5
x= -385/-5
x=77 ppm - original